Analytical Geometry

If P(x1,y1), Q(x2, y2),R(x3, y3)&S(x4,y4) are four concyclic points on the rectangular hyperbola xy=c^2, the coordinates of the orthocentre of the triangle PQR areAns(-x4,-y4)Plz give the full solution!!!!!

aayushi verma
8 Years agoGrade

1 Answer

Patel Dev Shaileshbhai
7 Years ago
H = xy = c2
(ct , c/t) is the harmonic coordinates 
 
from equation of circle 
 
x2 + y2 + 2gx + 2fy + c= 0
then 
 
c2t+ c2/t2 + 2gct + 2fc/t + c= 0
 
-t + 2gt3 + 2ft/c + c/t+ 1 = 0 
 
if tt2 tt4 on the circle 
then 
 
t1t2t3t= – 1 .
 
AND THEN ,
 
t= -t
 
so , ANS (-x4 ,-y4)